# Real Noncommutative Convexity II: Extremality and nc convex functions

### v1

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
[…]
@@
The following is the real case of Davidson and Kennedy’s noncommutative analogue of Milman’s partial converse to the Krein-Milman theorem [14, Theorem 6.4.3]:
@@
Let $K$ be a compact real nc convex set and $L$ be the noncommutative state space of $C^{*}_{\min}(A(K))$, identified with the set of noncommutative states on $C(K)$ which factor through $C^{*}_{\min}(A(K))$. Then, $L$ is the closed nc convex hull of $\{\delta_{x}\in\partial K\}$.
We can identify $L$ as the nc Shilov boundary of $V=A(K)$.
@@
(See [33] for an alternative take on some of the facts concerning $\mathbb{C}_{r}$ below.) Note that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. This follows from [29, Proposition 5.3.7 (1)] because these algebras have only one real state, hence only one pure real state. Thus e.g. all irreps of $\mathbb{H}$ are 4 dimensional and unitarily equivalent to the standard matrix form of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. Thus the real nc Choquet boundary of $\mathbb{C}_{r}$ is the set of elements $U^{T}\,c\,U$ where $c:\mathbb{C}\to M_{2}(\mathbb{R})$ is as usual, and $U$ is a rotation or reflection matrix. On the other hand, the complexification of $\mathbb{C}$ has two irreducible representations, and the complex nc Choquet boundary consists of two points. This shows that in general for a general nc convex set $K$, there need be no nice relation between the real Choquet boundary of $A(K)$ and the complex Choquet boundary in $K_{c}$. However the nc Shilov boundaries (which can be identified with the $C^{*}$-envelopes) are of course closely related. The noncommutative state space of $\mathbb{C}_{r}$ is identifiable with the nc convex set ${\mathcal{C}S}$ of skew symmetric real matrices of norm 1$\leq$ 1≤ 1. One can show by functional calculus that these are exactly the matrices $\varphi(i)$ for real ucp $\varphi:\mathbb{C}\to M_{n}(\mathbb{R})$. The noncommutative state space of $\mathbb{H}$ was discussed earlier and in [8]. These two sets coincide with the nc Shilov boundary, and are the nc closed convex hull of the nc Choquet boundary above, by Theorem 2.21. For the $\mathbb{C}$ example this reveals an interesting relation between the skew symmetric real matrices of norm 1$\leq$ 1≤ 1, and the single matrix $c(i)=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}.$ Namely, they are the noncommutative convex hull of $c(i)$: that is, ${\mathcal{C}S}={\rm nconv}(c(i))$. All of the $C^{*}$-extreme points of this set are computed in [33], and these turn out to correspond to the *-representations. So they are exactly the maximal elements by Proposition 2.12. The nc extreme points will be those *-representations which are irreducible, and we said that these are unitarily equivalent to the one in Example 3.5. The complexification of the nc convex set ${\mathcal{C}S}$ turns out to be the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. See Example 3.6 for more details. It would be quite interesting to perform a similar analysis to the above with the quaternions.
@@
The intersection in $C^{*}_{\rm min}(V)$ of the kernels of the boundary representations is (0)(0)( 0 ) (e.g. see the discussion above Theorem 2.15). Hence the corollary follows by the real case of basic facts about the topology of the spectrum of a $C^{*}$-algebra. Some of these basic facts may be found in [29] (e.g. Section 5.3 there), and in places in the proof of Theorem 5.3 in [9]). Other aspects of the real case of classical results about the spectrum $\hat{A}$ in early chapters of [17] need checking. We leave this to the interested reader. ∎
@@ 3. Complexification of extremals
We shall see later that a (scalar valued) state $\varphi$ on a real operator system $V$ is nc extreme if and only if $\varphi_{c}$ is complex nc extreme on $V_{c}$. In the language above, for every pure element $x\in K_{1}$ there exists $y$ with $c(x,y)\in K_{2}$ and $x+iy$ pure in $K_{c}$ (or equivalently, $c(x,y)$ pure in $K_{2}$).
@@
For $K$ a real compact nc convex set, if $x\in K$ is pure (resp. extreme) in $K_{c}$ then $x$ is pure (resp. extreme) in $K$. The converse is false: pure in $K$ need not imply pure (resp. extreme) in $K_{c}$. In particular, in general
@@
The converses of these are false even if $K$ corresponds to a commutative real $C^{*}$-algebra [33]. For example on $A=\{f:\mathbb{T}\to\mathbb{C}:f(\bar{z})=\overline{f(z)}\}$ the state $f\mapsto{\rm Re}\,f(i)$ is classically pure and extreme in $S(A)$, thus is nc pure, but is not pure nor extreme in any reasonable sense in $K_{c}$. Also consider the trace on $M_{2}(\mathbb{C})$. See also Example 3.5. ∎
@@
[…]
@@
Remark. For an element $x$ of a general complex compact nc convex set $K$, we do not know if $x$ is complex maximal in $K$ if it is real maximal in $K$. However we just proved that this does hold if $K$ is a complexification, and the next proposition is a partial result in this direction. It is worth remarking that any complex compact nc convex set that is not a complexification, is probably in a sense not easy to describe. Indeed such would give a complex operator system that is not an operator system complexification, and there are no known simple examples of this at this date (we thank Chris Phillips for a recent conversation on this point). For $C^{*}$-algebras this is equivalent to the hard problem of the existence of a real form (see e.g. [37]). There are not even any known simple low dimensional examples of complex Banach spaces that are not complexifications.
@@
. Suppose that $\varphi=\alpha^{*}\psi\alpha$ for $\psi\in{\rm UCP}_{\mathbb{C}}(B,M_{m}(\mathbb{C}))$ and real isometry $\alpha\in M_{m,n}(\mathbb{R})\subset M_{m,n}(\mathbb{C})$. View $\psi$ as in ${\rm UCP}_{\mathbb{R}}(B,M_{2m}(\mathbb{R}))$ and $c_{n}\circ\varphi$ in ${\rm UCP}_{\mathbb{R}}(B,M_{2n}(\mathbb{R}))$. Then $c_{n}\circ\varphi=c(\alpha+i0)^{\mathsf{T}}\,c_{m}\circ\psi\,c(\alpha+i0)$. Since $c_{n}\circ\pi$ is real maximal in $\mathrm{ncS}_{\mathbb{R}}(B)$ we have that $c(\alpha+i0)c(\alpha+i0)\mathsf{T}=c(\alpha\alpha^{*}+i0)$ commutes with $c_{m}\circ\psi$. However this forces $\alpha\alpha^{*}$ to commute with $\psi$ as desired. ∎
@@
$K=\mathrm{ncconv}\{c(i),c(-i)\}=\mathrm{ncconv}\{\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}\}\subset\mathcal{M}(\mathbb{R}).$
From Theorem 3.13.13.1 in [8], we know that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\mathrm{ncconv}_{\mathbb{R}}\{c(i)\}\}=\mathrm{ncconv}_{\mathbb{C}}\{c(i)\}$. A compression by $u$ shows ±$\pm i\in K_{c}$ so that $\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}\subseteq K_{c}$. Also,
@@
so that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}$. Let $L=\mathrm{ncconv}_{\mathbb{C}}\{\pm 1\}$, the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. Because ±$\pm i\in i\cdot L$ we have that $K_{c}\subseteq iL$. Conversely, ±$\pm 1\in-i\cdot\mathrm{ncconv}\{\pm i\}$ so that $L\subseteq-iK_{c}$. This implies $iL\subseteq K_{c}$ and thus $iL=K_{c}$. However, $L=\mathrm{ncconv}\{\pm 1\}=\bigsqcup_{n}[-1,1]_{n}$ by Example 4.14.14.1 in [40]. Also, $iL\cong L$ because multiplication by $i$ is a nc affine homeomorphism. Now, $L$ is compact, implying that $K_{c}$ is too, and thus $K$ is compact by Theorem 3.1. We have that ±$\pm 1\in L$ are the only nc extreme points also by Example 4.14.14.1 in [40], so that ±$\pm c(i)$ are the only possible extreme points in $K$, and they are indeed extreme. In Example 2.22 we observed that $K=\mathrm{ncS}(\mathbb{C}_{r})$, and are the skew matrix set ${\mathcal{C}S}$. So $A(K)\cong\mathbb{C}_{\mathbb{R}}$ as operator systems.
@@
[…]
@@
Because $\theta$ is bijective, this is a *-isomorphism. Although both $I$ and ${\mathcal{C}S}=K$ have the same complexification they are not nc affine homeomorphic, indeed their associated operator systems are $\ell^{\infty}_{2}(\mathbb{R})$ and $\mathbb{C}_{r}$, which are quite different. Indeed although $I={\rm Min}([-1,1])$ in the language of [8, Section 6], $K$ is not a Min nc convex set. Indeed $A(K)=\mathbb{C}_{r}$ is not an real OMIN operator system: it has nontrivial involution and is not a real function system in the technical sense of [9, 8].
@@
The converse of the lemma is false even in the simplest example of $K$ singleton: 111 is pure and extreme in $\mathrm{ncS}(\mathbb{C})$ but $c(1)=1\oplus 1$ is not pure, and is clearly not extreme. The class of extreme points considered in this lemma is interesting in some examples, but not in others.
@@
2) (A quaternionic counterexample.) Similarly the identity map on $M_{2}(\mathbb{C})$ is extreme, so maximal. However its real part $\psi$, which is Re, restricted to the copy of the quaternions $\mathbb{H}$, is not real maximal, so not extreme. This can be seen from Proposition 2.12 (2), or by the following direct argument: $\psi$ is the compression to the first two rows and columns of the usual (irreducible) representation of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. So it is not maximal by Proposition 2.12 (2). It is interesting that $\psi$ is pure, by the real case of Example 6.1.7 in [14]. (Cf. [33, Example 3.27].)
The latter also answers the question: For a real ucp map $\varphi$, there exists a maximal dilation of its complexification. Is the real part of this maximal dilation is a maximal dilation of $\varphi$? In the example above, the complexification of $\psi$ has maximal dilation $x\mapsto x\oplus x$ for $x\in M_{2}(\mathbb{C})$. The real part of this is a real dilation of $\psi$. It is the map $\psi\oplus\psi$. However this is not maximal, since a summand of a maximal is easily seen to be maximal, and $\psi$ is not maximal. This also follows by Proposition 2.12 (2).
@@
Remark. By a slight variant of the argument for the last assertion of Corollary 3.2, $x\in K_{1}$ is real extreme in $K_{1}$ if and only if $x$ is complex extreme in $K$. However generally the converses in Lemma 3.8 fail for pure (resp. extreme). E.g. if $x$ is real extreme in $K$ then $x$ need not be complex extreme in $K$. This fails even in $K_{2}$. To see this consider an extreme $x\in K$ such that $x$ is not complex nc extreme in $K_{c}$, such as in Example 3.5. However by Remark 2 after Example 3.5, $x$ is real nc extreme in $K_{c}$. Similarly $x$ is real pure in $K_{c}$, but not complex pure there (if it were it would be complex extreme since it is also complex maximal).

### v2

@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
As in [8, Section 6], $A(K)$ is an OMIN operator system unitally completely order isomorphic to OMIN(A)(A). Clearly the $C^{*}$-envelope $C^{*}_{\rm min}(A(K))$ is a real commutative $C^{*}$-algebra. However we need it to be the continuous functions on a compact set $X$, and this is not automatic in the real case. Indeed this follows here because it is a quotient of the real valued continuous functions on $C$ (we are using Stone-Weierstrass to see that $A$ generates the latter; also the quotient of a $C(X)$ space is a $C(X)$ space). As in the complex case explained there, or e.g. in 4.3.4 in [7], this implies that $X$ is the classical Shilov boundary of $A$, which is viewed as a closed subset of $C$. To see that $\partial K\subset K_{1}=C$, and $\partial K$ is the classical Choquet boundary in $C$, the argument in [14] works verbatim. For example, that if $x\in K_{1}$ is real extreme in $K$ then $x$ is an extreme point of the real state space of $A$, and hence is in the classical Choquet boundary. It is well known in convexity theory that the basic classical Choquet and Shilov boundary theory of a real function space works the same as the complex theory, except that signed measures are used rather than complex ones, via the real version of the Riesz representation theorem for measures. ∎
@@
The proof is the same as in the complex case, following quickly from Theorem 2.17 and [8, Theorem 3.6]. See also the real matrix Krein-Milman theorem in [22]. The following is the real case of Davidson and Kennedy’s noncommutative analogue of Milman’s partial converse to the Krein-Milman theorem [14, Theorem 6.4.3]:
@@
Let $K$ be a compact real nc convex set and let $D$ be the noncommutative state space of $C^{*}_{\min}(A(K))$, identified with the set of noncommutative states on $C(K)$ which factor through $C^{*}_{\min}(A(K))$. Then $D$ is the closed nc convex hull of $\{\delta_{x}\in\partial K\}$.
The proof is the same as in the complex case. We can identify $D$ as the nc Shilov boundary of $V=A(K)$.
@@
(See [33] for an alternative take on some of the facts concerning $\mathbb{C}_{r}$ below.) Note that $\mathbb{H},\mathbb{C}_{r}$ each have only one real *-irrep up to unitary equivalence. This follows from [29, Proposition 5.3.7 (1)] because these algebras have only one real state, hence only one pure real state. Thus e.g. all irreps of $\mathbb{H}$ are 4 dimensional and unitarily equivalent to the standard matrix form of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. Thus the real nc Choquet boundary of $\mathbb{C}_{r}$ is the set of elements $U^{T}\,c\,U$ where $c:\mathbb{C}\to M_{2}(\mathbb{R})$ is as usual, and $U$ is a rotation or reflection matrix. On the other hand, the complexification of $\mathbb{C}$ has two irreducible representations, and the complex nc Choquet boundary consists of two points. This shows that in general for a general nc convex set $K$, there need be no nice relation between the real Choquet boundary of $A(K)$ and the complex Choquet boundary in $K_{c}$. However the nc Shilov boundaries (which can be identified with the $C^{*}$-envelopes) are of course closely related. The noncommutative state space of $\mathbb{C}_{r}$ is identifiable with the nc convex set ${\mathcal{C}}{\mathcal{S}}$ of skew symmetric real matrices of norm $\leq 1$. One can show by functional calculus that these are exactly the matrices $\varphi(i)$ for real ucp $\varphi:\mathbb{C}\to M_{n}(\mathbb{R})$. The noncommutative state space of $\mathbb{H}$ was discussed earlier and in [8]. These two sets coincide with the nc Shilov boundary, and are the nc closed convex hull of the nc Choquet boundary above, by Theorem 2.21. For the $\mathbb{C}$ example this reveals an interesting relation between the skew symmetric real matrices of norm $\leq 1$, and the single matrix $c(i)=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$. Namely, they are the noncommutative convex hull of $c(i)$: that is, ${\mathcal{C}}{\mathcal{S}}={\rm nconv}(c(i))$. All of the $C^{*}$-extreme points of this set are computed in [33], and these turn out to correspond to the *-representations. So they are exactly the maximal elements by Proposition 2.12. The nc extreme points will be those *-representations which are irreducible, and we said that these are unitarily equivalent to the one in Example 3.5. The complexification of the nc convex set ${\mathcal{C}}{\mathcal{S}}$ turns out to be the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. See Example 3.6 for more details. It would be quite interesting to perform a similar analysis to the above with the quaternions.
@@
The intersection in $C^{*}_{\rm min}(V)$ of the kernels of the boundary representations is $(0)$ (e.g. see the discussion above Theorem 2.15). Hence the corollary follows by the real case of basic facts about the topology of the spectrum of a $C^{*}$-algebra. Some of these basic facts may be found in [29] (e.g. Section 5.3 there), and in places in the proof of Theorem 5.3 in [9]), or have already been discussed in the proof of Theorem 2.20 above. Any other aspects of the real case of classical results about the spectrum $\hat{A}$ in early chapters of [17] which we need here are easy to check and left to the interested reader. ∎
@@ 3. Complexification of extremals
We shall see later that a (scalar valued) state $\varphi$ on a real operator system $V$ is nc extreme if and only if $\varphi_{c}$ is complex nc extreme on $V_{c}$. Similarly we shall see that, in the language above, for every pure element $x\in K_{1}$ there exists $y$ with $c(x,y)\in K_{2}$ and $x+iy$ pure in $K_{c}$ (which follows if $c(x,y)$ is pure in $K_{2}$).
@@
For $K$ a real compact nc convex set, if $x\in K$ is pure (resp. extreme) in $K_{c}$ then $x$ is pure (resp. extreme) in $K$. The converse is false: pure (resp. extreme) in $K$ need not imply pure (resp. extreme) in $K_{c}$. In particular, in general
@@
The converses of these are false even if $K$ corresponds to a commutative real $C^{*}$-algebra [33]. For example on $A=\{f:\mathbb{T}\to\mathbb{C}:f(\bar{z})=\overline{f(z)}\}$ the state $f\mapsto{\rm Re}\,f(i)$ is classically pure and extreme in $S(A)$, thus is nc pure, but is not pure nor extreme in any reasonable sense in $K_{c}$. Also consider the restriction of the trace on $M_{2}(\mathbb{C})$ to the copy of the quaternions. See also Example 3.5. ∎
@@
From the last paragraph we have reduced checking that $x$ is extreme in $K_{c}$ to the special case: if $x=\sum_{k}\,a_{k}xa_{k}$ for positive $a_{k}\in M_{n}(\mathbb{C})$ with $\sum_{k}\,a_{k}^{2}=1$ and $a_{k}x=xa_{k}$, then $a_{k}$ is a positive scalar multiple of $I_{n}$. That is, for positive $b_{1},\cdots,b_{m}\in M_{n}(\mathbb{C})$ with $\sum_{k}\,b_{k}=1$ and $b_{k}x=xb_{k}$, then $b_{k}$ is a positive scalar multiple of $I_{n}$. It is easy to see that in this last condition we may assume that $m=2$. If this holds with $m=2$, and if $b$ is positive in $M_{n}(\mathbb{C})$ and $b$ commutes with $x$, then $\|b\|I-b$ satisfies the same conditions. So $b$ is a positive scalar multiple of $I_{n}$. Thus $x$ is extreme in $K_{c}$ if and only if every nonzero positive in $M_{n}(\mathbb{C})$ commuting with $x$ is a positive scalar multiple of $I_{n}$. Now the result is clear. Indeed the commutant in $M_{n}(\mathbb{C})$ of the range of $x$ is a von Neumann algebra, so is trivial if and only if it contains no nontrivial projections.
@@
Certainly every complex maximal is real maximal. For the other, suppose $x\in(K_{c})_{n}$ is real maximal. Write $x=\alpha^{*}y\alpha$ for $y\in(K_{c})_{l}$ and $\alpha\in M_{l,n}(\mathbb{C})$ an isometry. Taking $c$ of both sides gives $c(x)=c(\alpha)^{\mathsf{T}}c(y)c(\alpha)$. Compressing both sides by $v=[1_{n}\;0]^{\mathsf{T}}$ gives
@@
Remark. For an element $x$ of a general complex compact nc convex set $K$, we do not know if $x$ is complex maximal in $K$ if it is real maximal in $K$. However we just proved that this does hold if $K$ is a complexification, and the next proposition is a partial result in this direction. It is worth remarking that any complex compact nc convex set that is not a complexification, is probably in a sense not easy to describe. Indeed such would give a complex operator system that is not an operator system complexification, and there are no known simple examples of this at this date (we thank Chris Phillips for a recent conversation on this point). For $C^{*}$-algebras this is equivalent to the hard problem of the existence of a real form (see e.g. first page or two of [37]). There are not even any known simple low dimensional examples of complex Banach spaces that are not (isometric) complexifications.
@@
Suppose that $\varphi=\alpha^{*}\psi\alpha$ for $\psi\in{\rm UCP}_{\mathbb{C}}(B,M_{m}(\mathbb{C}))$ and real isometry $\alpha\in M_{m,n}(\mathbb{R})\subset M_{m,n}(\mathbb{C})$. View $\psi$ as in ${\rm UCP}_{\mathbb{R}}(B,M_{2m}(\mathbb{R}))$ and $c_{n}\circ\varphi$ in ${\rm UCP}_{\mathbb{R}}(B,M_{2n}(\mathbb{R}))$. Then $c_{n}\circ\varphi=c(\alpha+i0)^{\mathsf{T}}\,c_{m}\circ\psi\,c(\alpha+i0)$. Since $c_{n}\circ\pi$ is real maximal in $\mathrm{ncS}_{\mathbb{R}}(B)$ we have that $c(\alpha+i0)c(\alpha+i0)^{\mathsf{T}}=c(\alpha\alpha^{*}+i0)$ commutes with $c_{m}\circ\psi$. However this forces $\alpha\alpha^{*}$ to commute with $\psi$ as desired. ∎
@@
[…]
From Theorem $3.1$ in [8], we know that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\mathrm{ncconv}_{\mathbb{R}}\{c(i)\}\}=\mathrm{ncconv}_{\mathbb{C}}\{c(i)\}$. A compression by $u$ shows $\pm i\in K_{c}$ so that $\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}\subseteq K_{c}$. Also,
@@
so that $K_{c}=\mathrm{ncconv}_{\mathbb{C}}\{\pm i\}$. Let $L=\mathrm{ncconv}_{\mathbb{C}}\{\pm 1\}$, the operator interval $\{x\in{\mathcal{M}}(\mathbb{C})_{\rm sa}:-I\leq x\leq I\}$. Because $\pm i\in i\cdot L$ we have that $K_{c}\subseteq iL$. Conversely, $\pm 1\in-i\cdot\mathrm{ncconv}\{\pm i\}$ so that $L\subseteq-iK_{c}$. This implies $iL\subseteq K_{c}$ and thus $iL=K_{c}$. However, $L=\mathrm{ncconv}\{\pm 1\}=\bigsqcup_{n}[-1,1]_{n}$ by Example $4.1$ in [40]. Also, $iL\cong L$ because multiplication by $i$ is a nc affine homeomorphism. Now, $L$ is compact, implying that $K_{c}$ is too, and thus $K$ is compact by Theorem 3.1. We have that $\pm 1\in L$ are the only nc extreme points also by Example $4.1$ in [40], so that $\pm c(i)$ are the only possible extreme points in $K$, and they are indeed extreme. In Example 2.22 we observed that $K=\mathrm{ncS}(\mathbb{C}_{r})$, and equals the skew matrix set ${\mathcal{C}}{\mathcal{S}}$. So $A(K)\cong\mathbb{C}_{\mathbb{R}}$ as operator systems.
@@
Let $I=\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}$. This equals the compact real operator interval $\bigsqcup_{n}[-1,1]_{n}$. To see this, $\subseteq$ is obvious, and for the converse we observe that $\mathrm{ncconv}_{\mathbb{R}}\{\pm 1\}=\mathrm{ncconv}_{\mathbb{R}}\{[-1,1]\}$. Elements of the compact operator interval can all be written as the diagonalization $U^{\mathsf{T}}DU$ for $D$ the direct sum of elements in the interval [1,1][-1,1]. Thus, all elements in the operator interval are in $I$. So $I$ is the compact real operator interval. Elements of $A(I)$ are determined by where they send $\pm 1$. Therefore, $A(I)$ is a two dimensional real operator system, indeed is a real $C^{*}$-algebra *-isomorphic to $\ell^{\infty}_{2}(\mathbb{R})$. The isomorphism is the map $\theta:A(I)\to\ell^{\infty}_{2}(\mathbb{R})$ taking $f\in A(I)$ to (f(1),f(1))(f(-1),f(1)). Note that $A(I)$ has trivial involution, so all we need to check is that $\theta$ is multiplicative. However,
@@
Because $\theta$ is bijective, this is a *-isomorphism. Although both $I$ and ${\mathcal{C}}{\mathcal{S}}=K$ have the same complexification they are not nc affine homeomorphic, indeed their associated operator systems are $\ell^{\infty}_{2}(\mathbb{R})$ and $\mathbb{C}_{r}$, which are quite different. Indeed although $I={\rm Min}([-1,1])$ in the language of [8, Section 6], $K$ is not a Min nc convex set. Indeed $A(K)=\mathbb{C}_{r}$ is not a real OMIN operator system: it has nontrivial involution (has a skew element) and is not a real function system in the technical sense of [9, 8]. Indeed we showed in [9, Section 10] that an operator system cannot be an OMIN if it has a skew element, and as we mentioned earlier a real $C^{*}$-algebra with no skew element is a real $C(X)$.
@@
The converse of the lemma is false even in the simplest example of $K$ singleton: $1$ is pure and extreme in $\mathrm{ncS}(\mathbb{C})$ but $c(1)=1\oplus 1$ is not pure, and is clearly not extreme. The class of extreme points considered in this lemma is interesting in some examples, but not in others.
@@
2) (A quaternionic counterexample.) Similarly the identity map on $M_{2}(\mathbb{C})$ is extreme, so maximal. However its real part $\psi$, which is Re, restricted to the copy of the quaternions $\mathbb{H}$, is not real maximal, so not extreme. This can be seen from Proposition 2.12 (2), or by the following direct argument: $\psi$ is the compression to the first two rows and columns of the usual (irreducible) representation of $\mathbb{H}$ in $M_{4}(\mathbb{R})$. So it is not maximal by Proposition 2.12 (2). It is interesting that $\psi$ is pure, by Proposition 2.12 (1). (Cf. [33, Example 3.27].)
The latter also answers the question: For a real ucp map $\varphi$, there exists a maximal dilation of its complexification. Is the real part of this maximal dilation a maximal dilation of $\varphi$? In the example above, the complexification of $\psi$ has maximal dilation $x\mapsto x\oplus x$ for $x\in M_{2}(\mathbb{C})$. The real part of this is a real dilation of $\psi$. It is the map $\psi\oplus\psi$. However this is not maximal, since a summand of a maximal is easily seen to be maximal, and $\psi$ is not maximal. This also follows by Proposition 2.12 (2).
@@
Remark. By a slight variant of the argument for the last assertion of Corollary 3.2, $x\in K_{1}$ is real extreme in $K_{1}$ if and only if $x$ is complex extreme in $K$. However generally the converses in Lemma 3.8 fail for pure (resp. extreme). E.g. if $x$ is real extreme in $K$ then $x$ need not be complex extreme in $K$. This fails even in $K_{2}$. To see this consider an extreme $x\in K$ such that $x$ is not complex nc extreme in $K_{c}$, such as in Example 3.5. However by the Remark after Example 3.5, $x$ is real nc extreme in $K_{c}$. Similarly $x$ is real pure in $K_{c}$, but not complex pure there (if it were it would be complex extreme since it is also complex maximal).